OpenStudy (anonymous):

You attend an amusement park with your family. Your parents buy you an all-ride pass for $20, shown as f(x). Instead of getting a pass, your parents decide to pay $4 for each ride they take, shown as g(x). What function shows the correct combination of these two functions to represent the total cost to them of attending the amusement park that day, shown as h(x)? f(x) = 20x, g(x) = 4, h(x) = 20x + 4 f(x) = 20, g(x) = 4, h(x) = 4 + 20 f(x) = 20, g(x) = 4x, h(x) = 4x + 20 f(x) = 20x, g(x) = 4x, h(x) = 20x + 4x

OpenStudy (anonymous):

@gorv

OpenStudy (donblue22):

The answer would be C since the parents are paying $4 a ride (4x) and 20 total for you

OpenStudy (anonymous):

This is easy, the first one is an all out pass which won't have a variable. The second one is per ride, so it has a variable.

OpenStudy (anonymous):

@Donblue22 Don't give out answers.

OpenStudy (anonymous):

Thank you Mrs. Greenlegodude57

OpenStudy (donblue22):

I didnt give him the answer I actually explained it

OpenStudy (anonymous):

LOL @Solver

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

You're a girl, right? Since there's a flower on your pic.

OpenStudy (anonymous):

@Donblue22 It's still a direct answer if you explain it. @Solver can just take the answer and not look at the explanation. So you have to guide him to the answer.

OpenStudy (anonymous):

@Donblue22 You did..you said 'The answer would be C'

OpenStudy (anonymous):

READ THIS FROM THE MODERATOR!

OpenStudy (donblue22):

Just leave me alone Im having a bad morning and your not heloing im just trying to be helpful

OpenStudy (anonymous):

You're not being helpful.

OpenStudy (gorv):

all ride pass cost $20 it is fixed

OpenStudy (gorv):

it do not change ..you can take any number of ride

OpenStudy (gorv):

so it is constant

OpenStudy (gorv):

@greenlegodude57 gotcha buddy ??

OpenStudy (anonymous):

??

OpenStudy (gorv):

what i told

OpenStudy (gorv):

f(x)=20 constant

OpenStudy (gorv):

let u took x ride

OpenStudy (gorv):

and each cost $4

OpenStudy (gorv):

so x ride will cost = 4*x=4x

OpenStudy (gorv):

g(x)=4x

OpenStudy (gorv):

so h(x)=f(x)+g(x) plug in what we got for f(x) and g(x)

OpenStudy (anonymous):

You do know that this is @Solver 's question, right? @gorv

OpenStudy (gorv):

my mistake @greenlegodude57